OpenStudy (anonymous):

could someone show me how to find the absolute values of |2x+3|=|3x-4|

OpenStudy (ash2326):

@jpjones do you know what does this imply |x|?

OpenStudy (anonymous):

absolute value

OpenStudy (ash2326):

good:) we have \[|2x+3|=|3x+4|\] there are 2 cases I) 2x+3>0 2) 2x+3<0

OpenStudy (anonymous):

is it better to interprate |x| as +x or -x

OpenStudy (ash2326):

yeah we'll get to that:) if 2x+3\(\ge\)0 then \[|2x+3|=2x+3\] do you agree?

OpenStudy (anonymous):

yes i agree

OpenStudy (ash2326):

let's find the ranges, I did a mistake somewhere

OpenStudy (ash2326):

if we solve \[2x+3\ge0\] we get \[x\ge \frac{-3}{2}\] so for \[x\ge -\frac{3}{2}=> |2x+3|=2x+3\] and for \[x< -\frac{3}{2}=> |2x+3|=-2x+3\] Now would you do the same for 3x+4 find the ranges of x

OpenStudy (anonymous):

i can tell you that the answer in the back of the book say's 1/5, and 7, so somewhere there is a mistake in your working, i must also explain that the question request all real values for x

OpenStudy (ash2326):

@jpjones we haven't solved it yet

OpenStudy (ash2326):

these are some points, which will help us to solve

OpenStudy (anonymous):

ok

OpenStudy (ash2326):

We'll start with a case let x be less than -3/2 then we'll get the equation as 3x+4<0 for x=-3/2 \[-(2x+3)=-(3x+4) \] now we'll solve it to find x if it comes out to be less than -3/2 then only it'll be an answer

OpenStudy (ash2326):

do you get my point?

OpenStudy (anonymous):

yes this statement is true but how do we solve from here

OpenStudy (ash2326):

@jpjones Sorry there is an easy way, do you know quadratic equations?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

just square both sides then \[(2x+3)^2=(3x-4)^2\] what would you get now?

OpenStudy (anonymous):

i'm a long way behind in my math and am trying to catch up, if you can't solve this let me know and i'll move on

OpenStudy (ash2326):

No, we can solve this:) would you simplify it?

OpenStudy (ash2326):

Sorry the other way was long but it was also correct

OpenStudy (anonymous):

i got 4x/9 =9x/16 cancel x, cancel 9 then left with 4/16 which is 1/4

OpenStudy (ash2326):

after you square it, you'd get \[4x^2+9+12x=9x^2+16-24x\] can you solve it now?

OpenStudy (anonymous):

no show me

OpenStudy (ash2326):

do you know the quadratic formula?

OpenStudy (ash2326):

@jpjones ??